There are also easy ways to draw commutative diagrams, like:

ℝ^{2} ⟶^{F} ℝ^{2}
h↓  ↓k
ℝ^{2} ⟶_{G} ℝ^{2}
ℝ^{2} ⟶_{F} ℝ^{2} ⟶_{H} ℝ^{2} ⟶^{I} ℝ^{2}
↓h  k↓
ℝ^{2} ⟶^{G} ℝ^{2}

Notice the different positions of the labels.